A random sample of 328 medical doctors showed that 172 had a solo practice.

(a) Let p represent the proportion of all medical doctors who have a solo practice. Find a point estimate for p. (Use 3 decimal places.)


(b) Find a 90% confidence interval for p. (Use 3 decimal places.)

lower limit
upper limit

Give a brief explanation of the meaning of the interval.

90% of the confidence intervals created using this method would include the true proportion of physicians with solo practices.

10% of the confidence intervals created using this method would include the true proportion of physicians with solo practices.  

  90% of the all confidence intervals would include the true proportion of physicians with solo practices.

10% of the all confidence intervals would include the true proportion of physicians with solo practices.


(c) As a news writer, how would you report the survey results regarding the percentage of medical doctors in solo practice?

Report the confidence interval.

Report .    

Report along with the margin of error.

Report the margin of error.



What is the margin of error based on a 90% confidence interval? (Use 3 decimal places.)
Answer #1

Solution :

Given that,

n = 328

x = 172

Point estimate = sample proportion = = x / n = 172/328=0.5244

1 - = 1-0.5244=0.4756

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645 ( Using z table )

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 1.645 (((0.5244*0.4756) /328 )

E = 0.045

A 90% confidence interval for population proportion p is ,

- E < p < + E

0.5244 - 0.045< p <0.5244 + 0.045

0.479< p < 0.569

The 90% confidence interval for the population proportion p is : lower limit =0.479 , upper limit 0.569

brief explanation =90% of the all confidence intervals would include the true proportion of physicians with solo practices.